# Bypassing the lattice BCS-BEC crossover in strongly correlated superconductors: resilient coherence from multiorbital physics

### v2

@@ Abstract
Superconductivity emerges from the spatial coherence of a macroscopic condensate of Cooper pairs. Increasingly strong binding and localization of electrons into these pairs compromises the condensate’s phase stiffness, thereby limiting critical temperatures – a phenomenon known as the BCS-BEC crossover in lattice systems. In this study, we report on an enhancement of superconductivity beyond the limits of the lattice BCS-BEC crossover realized in a multiorbital model of alkali-doped fullerides (A3C60). We show how strong correlations and multiorbital effects lead into a localized superconducting regime characterized by a short coherence length but robust stiffness and a domeless rise in critical temperature with increasing pairing interaction. These insights are derived from the development of a theoretical framework to calculate the fundamental length scales of superconductors, namely the coherence length ($\xi_{0}$) and the London penetration depth ($\lambda_{\mathrm{L}}$), in microscopic theories and from first principles, even in presence of strong electron correlations.
coherence length, correlation length, penetration depth, fulleride, A3C60, alkali, doped, Buckyball, Buckminster, superconductivity, superconductor, fluctuations, s-wave, order parameter, depairing current, critical current, supercurrent, critical fields, finite-momentum pairing, helical phase, FFLO, Hund’s metal, strong coupling, multiorbital, Ginzburg-Landau, type-I, type-II, strong correlation, phase gradient, phase twist, stiffness, DMFT, Dynamical Mean-Field Theory, BCS to BEC crossover, BEC-BCS crossover, BEC, BCS, Uemura plot, superconducting dome
@@ Introduction
The collective and phase coherent condensation of electrons into bound Cooper pairs leads to the emergence of superconductivity. This macroscopic coherence enables dissipationless charge currents, perfect diamagnetism, fluxoid quantization and technical applications (1, 2) ranging from electromagnets in particle accelerators to quantum computing hardware. Often, superconducting (SC) functionality is controlled by the critical surface spanned by critical magnetic fields, currents, and temperatures which a SC condensate can tolerate. Fundamentally, these are determined by the characteristic length scales of a superconductor – the London penetration depth, $\lambda_{\mathrm{L}}$, and the coherence length, $\xi_{0}$.
@@
$\lambda_{\mathrm{L}}$ and $\xi_{0}$ quantify different aspects of the SC condensate: The penetration depth is the length associated with the mass term that the vector potential gains through the Anderson-Higgs mechanism (3). In consequence, magnetic fields decay exponentially over a distance $\lambda_{\mathrm{L}}$ inside a superconductor. Through this, $\lambda_{\mathrm{L}}$ is connected to the energy cost of order parameter (OP) phase variations and hence the SC stiffness $D_{\mathrm{s}}$. The coherence length, on the other hand, is the intrinsic length scale of OP amplitude variations and is associated with the amplitude Higgs mode. $\xi_{0}$ sets the scale below which amplitude and phase modes significantly couple such that spatial variations of the OP’s phase reduce its amplitude (3, 4).
@@
In addition to influencing the macroscopic properties of superconductors, $\lambda_{\mathrm{L}}$ and $\xi_{0}$ play an important role to understand strongly correlated superconductors, as is epitomized in the Uemura plot (5, 6, 7, 8). For instance, the interplay of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ impacts critical temperatures (9), it is relevant for the pseudogap formation (10, 11, 12, 13), and it might underlie the light-enhancement of superconductivity (14, 15, 16, 17, 18). An important concept in this context is the BCS-BEC crossover phenomenology (19, 20, 21). It continuously connects the two limiting cases of weak-coupling Bardeen-Cooper-Schrieffer (BCS) superconductivity with weakly-bound and largely overlapping Cooper pairs to tightly-bound molecule-like pairs in the strong-coupling Bose-Einstein condensate (BEC) as the interaction strength or the density is varied (Fig. 1).
@@
The BCS-BEC crossover has been studied in ultracold Fermi gases (20), low-density doped semiconductors (22), and is under debate for several unconventional superconductors (6, 8, 19, 23, 24, 25, 26). However, quasi-continuous systems, for instance Fermi gases, and strongly correlated superconducting solids show a crucially different behavior of how their SC properties, most importantly the critical temperature $T_{\mathrm{c}}$, change towards the strong coupling BEC limit. While $T_{\mathrm{c}}$ converges to a constant temperature TBEC for Fermi gases in a continuum (Fig. 1A), $T_{\mathrm{c}}$ can become arbitrarily small in strongly correlated lattice systems due to the quenching of kinetic energy (Fig. 1B). Since the movement of electron pairs, i.e., bosonic hopping, necessitates intermediate fermionic hopping, it becomes increasingly unfavorable for strong attractions. Thus, as Cooper pairs become localized on the scale of the lattice constant, the condensate’s stiffness and hence $T_{\mathrm{c}}$ are compromised (19, 21). Fig. 1 contrasts this generic BCS-BEC crossover picture for Fermi gases and correlated lattice systems in terms of the change of $T_{\mathrm{c}}$ and the pair size $\xi$ as a function of pairing strength. Due to the decrease of $T_{\mathrm{c}}$ in the BCS and BEC limits, a prominent dome-shape of $T_{\mathrm{c}}$ can be expected in the crossover region for solid materials. Because of this, recent experimental efforts to increase $T_{\mathrm{c}}$ concentrate on stabilizing SC materials in this region (19, 22, 24, 26).
@@
Fig. 1: BCS-BEC crossover in Fermi gases vs. lattice systems. Evolution of the critical temperature $T_{\mathrm{c}}$ and Cooper pair size characterized by the coherence length $\xi$ in the BCS-BEC crossover (dark blue line) for (A) Fermi gases and (B) lattice systems. Both cases display a dome-shaped behavior of $T_{\mathrm{c}}$ in the intermediate crossover regime but behave qualitatively different in the strong coupling BEC phase: $T_{\mathrm{c}}$ remains finite in the Fermi gas (A) but approaches zero in the lattice case (B). During the crossover, $\xi$ is reduced to the order of interparticle spacing $r$ (A & B) and lattice constant $a$ (B). For the lattice system, we contrast the evolution towards the multiorbital strong coupling phase (light blue line) discussed in this article. Here, the localization of pairs differently affects the Bosonic hopping $t_{\mathrm{B}}$, as drawn in the insets on the right. In the lattice BEC limit, Bosonic hopping relies on a second-order process involving two fermionic hoppings, $t_{\mathrm{F}}$, and a virtual intermediate state with broken pairs that is inhibited by the strong attraction $g<0$. Consequently, this process and also $T_{\mathrm{c}}\propto t_{\mathrm{B}}=t_{\mathrm{F}}^{2}/|g|$ are quenched at large $|g|$ (19, 21). In the multiorbital strong coupling case, the local coexistence of paired and unpaired electrons fluctuating between different orbitals enables Bosonic hopping $t_{\mathrm{B}}$ as a first order process in $t_{\mathrm{F}}$ without any intermediate broken pair states. A second temperature scale $T^{*}$ is drawn in both panels as its splitting from $T_{\mathrm{c}}$ (corresponding to the opening of a pseudogap) marks the beginning of the crossover regime (19).
@@
In this work, we demonstrate how multiorbital effects can enhance superconductivity beyond the expectations of the lattice BCS-BEC crossover phenomenology (as contrasted in Fig. 1B) with a model inspired by alkali-doped fullerides (A3C60 with A=K, Rb, Cs). The material family of A3C60 hosts exotic $s$-wave superconductivity of critical temperatures up to $T_{\mathrm{c}}=38$ K, being the highest temperatures among molecular superconductors (27, 28, 29, 30), and they possibly reach photo-induced SC at even higher temperatures (14, 16, 17, 31). In order to theoretically characterize the SC state, the knowledge of the intrinsic SC length scales is essential. While BCS theory and Eliashberg theory provide a microscopic description of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ for weakly correlated materials (32, 33, 34), their validity is unclear for superconductors with strong electron correlations. To the best of our knowledge, $\xi_{0}$ is generally not known from theory in strongly correlated materials. Only approaches to determine $\lambda_{\mathrm{L}}$ exist where an approximate, microscopic assessment of the SC stiffness from locally exact theories has been established, albeit neglecting vertex corrections (35, 36, 37, 38, 39).
@@
Here, we introduce a novel theoretical framework to microscopically access $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from the tolerance of SC pairing to spatial OP variations. Central to our approach are calculations in the superconducting state under a constraint of finite-momentum pairing (FMP). Via Nambu-Gor’kov Green functions we get direct access to the superconducting OP and the depairing current $j_{\mathrm{dp}}$ which in turn yield $\xi_{0}$ and $\lambda_{\mathrm{L}}$. The FMP constraint is the SC analog to planar spin spirals applied to magnetic systems (40, 41, 42). As in magnetism, a generalized Bloch theorem holds that allows us to consider FMP without supercells; see section S2 in the supplementary text for a proof. As a result, our approach can be easily embedded in microscopic theories and ab initio approaches to tackle material-realistic calculations. In this work, we implement FMP in Dynamical Mean-Field theory (DMFT) (43) to treat strongly correlated superconductivity. In DMFT, the interacting many-body problem is solved by self-consistently mapping the lattice model onto a local impurity problem. By this, local correlations are treated exactly.
@@
We apply the FMP-constrained in DMFT to A3C60 and we find $\xi_{0}$ and $\lambda_{\mathrm{L}}$ in line with experiment for model parameter ranges derived from ab initio estimates, validating our approach. For enhanced pairing interaction, we then reveal a multiorbital strong coupling SC state with minimal $\xi_{0}$ on the order of only 2 – 3 lattice constants, but with robust stiffness $D_{\mathrm{s}}$ and high $T_{\mathrm{c}}$ which increases with the pair interaction strength without a dome shape. This strong coupling SC state circumvents the lattice BCS-BEC crossover phenomenology. In the following, we motivate how the FMP constraint is linked to $\lambda_{\mathrm{L}}$ and $\xi_{0}$ with the subsequent application to the multiorbital model of A3C60.
@@ Results
Superconductivity with the constraint of finite-momentum pairing
@@
In most SC materials, Cooper pairs do not carry a finite center-of-mass momentum $\bm{q}=\bm{0}$. Yet, in presence of external fields, coexisting magnetism, or even spontaneously SC states with FMP, i.e., $\bm{q}\neq\bm{0}$, might arise (44, 45, 46, 47, 48) as originally conjectured in Fulde-Ferrel-Larkin-Ovchinnikov (FFLO) theory (49, 50, 51). Here, we enforce FMP states via constraints on the pair-center-of mass momentum $\bm{q}$ in order to access the characteristic SC length scales, $\xi_{0}$ and $\lambda_{\mathrm{L}}$, through an analysis of the momentum- and temperature-dependent OP and the closely connected charge supercurrent. We constrain the OP to be of the FMP form $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ corresponding to FF-type pairing (49), which is to be differentiated from pair density waves with amplitude modulations (50, 51, 52). We contrast the OP for zero and finite momentum in real and momentum space in the top panel of Fig. 2. For FMP, the OP’s phase is a helix winding around the direction of $\bm{q}$, while the OP for zero momentum pairing is simply a constant.
@@
Before turning to the implementation in microscopic approaches, we motivate how the FMP constraint relates to $\lambda_{\mathrm{L}}$ and $\xi_{0}$. The Ginzburg-Landau (GL) framework provides an intuitive picture to this connection which we summarize here and discuss in detail in section S1 of the supplementary text. The GL low-order expansion of the free energy density $f_{\mathrm{GL}}$ in terms of the FMP-constrained OP reads
$f_{\mathrm{GL}}[\Psi_{\bm{q}}]=a|\Psi_{\bm{q}}|^{2}+\frac{b}{2}|\Psi_{\bm{q}}|^{4}+\frac{\hbar^{2}}{2m^{*}}q^{2}|\Psi_{\bm{q}}|^{2}$ (1)
@@
with $q=|\bm{q}|$. $a$, $b$, and $m^{*}$ are the material and temperature dependent GL parameters. Here, the temperature dependent (GL) correlation length $\xi$ appears as the natural length scale of the amplitude mode ($\propto a$) and the kinetic energy term
$\xi(T)=\sqrt{\frac{\hbar^{2}}{2m^{*}|a|}}=\xi_{0}\left(1-\frac{T}{T_{\mathrm{c}}}\right)^{-\frac{1}{2}}$ (2)
@@
with its zero-temperature value $\xi_{0}$ being the coherence length (4). The stationary point of Eq. (1) shows that the $\bm{q}$-dependent OP amplitude $\left|\Psi_{q}\right|=\left|\Psi_{0}\right|\sqrt{1-\xi^{2}q^{2}}$ decreases with increasing momentum $q$. For large enough $q$, SC order breaks down (i.e., $\Psi_{q}\to 0$) as the kinetic energy from phase modulations becomes comparable to the gain in energy from pairing. The length scale associated with this breakdown is $\xi$ and can, therefore, be inferred from the $q$-dependent OP suppression. We employ, here, the criterion $\xi=1/(\sqrt{2}|\bm{Q}|)$ with $\bm{Q}$ such that $|\Psi_{\bm{Q}}/\Psi_{0}|=1/\sqrt{2}$ (see section S5.A of the supplementary text for more information).
@@
The finite center-of-mass momentum of the Cooper pairs entails a charge supercurrent $\bm{j}_{\bm{q}}\propto|\Psi_{\bm{q}}|^{2}\bm{q}$, c.f. Eq. (S18) in the supplementary text. This current density is a non-monotonous function of $q$ with a maximum called depairing current density $j_{\mathrm{dp}}$. It provides a theoretical upper bound to critical current densities, $j_{\mathrm{c}}$, measured in experiment. We note that careful design of SC samples is necessary for $j_{\mathrm{c}}$ reaching $j_{\mathrm{dp}}$ as its value crucially depends on sample geometry and defect densities (33, 53). $j_{\mathrm{dp}}$ is related to the London penetration depth via
$\lambda_{\mathrm{L}}(T)=\sqrt{\frac{\Phi_{0}}{3\sqrt{3}\pi\mu_{0}\xi j_{\mathrm{dp}}}}=\lambda_{\mathrm{L},0}\left(1-\left(\frac{T}{T_{\mathrm{c}}}\right)^{4}\right)^{-\frac{1}{2}}$ (3)
@@
with the magnetic flux quantum $\Phi_{0}=h/2e$. The temperature dependence with the quartic power stated here is empirical (32, 33) and we find that it describes our calculations better than the linearized GL expectation as discussed in the supplementary text (section S5.B).
@@
Fig. 2: Influence of finite-momentum pairing (FMP) constraint on the superconducting condensate. The top panel insets sketch the position and momentum space representation of the order parameter (OP) $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ in the zero-momentum (left, $q=0$) and finite-momentum pairing states (right, $q>0$). The main panels show (A) the momentum dependence of the OP modulus and (B) the supercurrent density $j_{q}=|\bm{j}_{\bm{q}}|$ as function of Cooper pair momentum $q=|\bm{q}|$ in reciprocal lattice units (r.l.u.). Gray lines indicate the points of extracting $\xi$ and $j_{\mathrm{dp}}$ (see text). The data shown are results for the A3C60 model (c.f. Eq. (7)) with interaction parameters $U/W=1.4$, $J/W=-0.04$ evaluated at different temperatures $T$ (color coded; see color bar with white triangular markers).
@@
The GL analysis shows that the OP suppression and supercurrent induced by the FMP constraint connect to $\xi$ and $\lambda_{\mathrm{L}}$. In a microscopic description, we acquire the OP and supercurrent density from the Nambu-Gor’kov Green function $\mathcal{G}$
@@
where $\psi_{\bm{k},\bm{q}}^{\dagger}=\left(c_{\bm{k}+\frac{\bm{q}}{2}\uparrow}^{\dagger}\ \ c_{-\bm{k}+\frac{\bm{q}}{2}\downarrow}\right)$ (orbital indices suppressed) are Nambu spinors that carry an additional dependence on $\bm{q}$ due to the FMP constraint. $G$ ($F$) denotes the normal (anomalous) Green function component for electrons ($G$) and holes ($\bar{G}$) in imaginary time $\tau$. For $s$-wave superconductivity as in the A3C60 model, we use the local anomalous Green function as the OP
@@
which is the same for all orbitals $\alpha$. The current density can be calculated via (c.f. Eq. (S37) in the supplementary text)

### v4

# Bypassing the lattice BCS–BEC crossover in strongly correlated superconductors through multiorbital physics

@@ Abstract
Superconductivity emerges from the spatial coherence of a macroscopic condensate of Cooper pairs. Increasingly strong binding and localization of electrons into these pairs compromises the condensate’s phase stiffness, thereby limiting critical temperatures – a phenomenon known as the BCS–BEC crossover in lattice systems. In this study, we demonstrate enhanced superconductivity in a multiorbital model of alkali-doped fullerides (A3C60) that goes beyond the limits of the lattice BCS–BEC crossover. We identify that the interplay of strong correlations and multiorbital effects results in a localized superconducting state characterized by a short coherence length but robust stiffness and a domeless rise in critical temperature with increasing pairing interaction. To derive these insights, we introduce a new theoretical framework allowing us to calculate the fundamental length scales of superconductors, namely the coherence length ($\xi_{0}$) and the London penetration depth ($\lambda_{\mathrm{L}}$), even in presence of strong electron correlations.
coherence length, correlation length, penetration depth, fulleride, A3C60, alkali, doped, Buckyball, Buckminster, superconductivity, superconductor, fluctuations, s-wave, order parameter, depairing current, critical current, supercurrent, critical fields, finite-momentum pairing, helical phase, FFLO, Hund’s metal, strong coupling, multiorbital, Ginzburg–Landau, type-I, type-II, strong correlation, phase gradient, phase twist, stiffness, DMFT, Dynamical Mean-Field Theory, BCS to BEC crossover, BEC-BCS crossover, BEC, BCS, Uemura plot, superconducting dome
@@ Introduction
The collective and phase coherent condensation of electrons into bound Cooper pairs leads to the emergence of superconductivity. This macroscopic coherence enables dissipationless charge currents, perfect diamagnetism, fluxoid quantization and technical applications 1, 2 ranging from electromagnets in particle accelerators to quantum computing hardware. Often, superconducting (SC) functionality is controlled by the critical surface spanned by critical magnetic fields, currents, and temperatures which a SC condensate can tolerate. Fundamentally, these are determined by the characteristic length scales of a superconductor – the London penetration depth, $\lambda_{\mathrm{L}}$, and the coherence length, $\xi_{0}$.
@@
$\lambda_{\mathrm{L}}$ and $\xi_{0}$ quantify different aspects of the SC condensate: The penetration depth is the length associated with the mass term that the vector potential gains through the Anderson-Higgs mechanism 3. In consequence, magnetic fields decay exponentially over a distance $\lambda_{\mathrm{L}}$ inside a superconductor. Through this, $\lambda_{\mathrm{L}}$ is connected to the energy cost of order parameter (OP) phase variations and hence the SC stiffness $D_{\mathrm{s}}$. The coherence length, on the other hand, is the intrinsic length scale of OP amplitude variations and is associated with the amplitude Higgs mode. $\xi_{0}$ sets the scale below which amplitude and phase modes significantly couple such that spatial variations of the OP’s phase reduce its amplitude 3, 4.
@@
In addition to influencing the macroscopic properties of superconductors, $\lambda_{\mathrm{L}}$ and $\xi_{0}$ play an important role to understand strongly correlated superconductors, as is epitomized in the Uemura plot 5, 6, 7, 8. For instance, the interplay of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ impacts critical temperatures 9, it is relevant for the pseudogap formation 10, 11, 12, 13, it influences magneto-thermal transport properties like the Nernst effect 8, 14, 15, and it might underlie the light-enhancement of superconductivity 15, 16, 17, 18, 19. An important concept in this context is the BCS–BEC crossover phenomenology 20, 21, 22, 23, 24. It continuously connects the two limiting cases of weak-coupling Bardeen–Cooper–Schrieffer (BCS) superconductivity with weakly-bound and largely overlapping Cooper pairs to tightly-bound molecule-like pairs in the strong-coupling Bose–Einstein condensate (BEC) as the interaction strength or the density is varied (Fig. 1).
@@
The BCS–BEC crossover has been studied in ultracold Fermi gases 23, low-density doped semiconductors 25, and is under debate for several unconventional superconductors 6, 8, 24, 26, 27, 28, 29, 30. However, quasi-continuous systems, for instance Fermi gases, and strongly correlated superconducting solids show a crucially different behavior of how their SC properties, most importantly the critical temperature $T_{\mathrm{c}}$, change towards the strong coupling BEC limit. While $T_{\mathrm{c}}$ converges to a constant temperature $T_{\mathrm{BEC}}$ for Fermi gases in a continuum 31 (Fig. 1a), $T_{\mathrm{c}}$ can become arbitrarily small in strongly correlated lattice systems due to the quenching of kinetic energy (Fig. 1b). Since the movement of electron pairs, i.e., bosonic hopping, necessitates intermediate fermionic hopping, it becomes increasingly unfavorable for strong attractions. Thus, as Cooper pairs become localized on the scale of the lattice constant, the condensate’s stiffness and hence $T_{\mathrm{c}}$ are compromised 24, 21. Fig. 1 contrasts this generic BCS–BEC crossover picture for Fermi gases and correlated lattice systems in terms of the change of $T_{\mathrm{c}}$ and the pair size $\xi_{\mathrm{p}}$ as a function of pairing strength. Due to the decrease of $T_{\mathrm{c}}$ in the BCS and BEC limits, a prominent dome-shape of $T_{\mathrm{c}}$ can be expected in the crossover region for solid materials. Because of this, recent experimental efforts to increase $T_{\mathrm{c}}$ concentrate on stabilizing SC materials in this region 24, 25, 27, 30.
@@
Fig. 1: BCS–BEC crossover in Fermi gases vs. lattice systems. Evolution of the critical temperature $T_{\mathrm{c}}$ and Cooper pair size $\xi_{\mathrm{p}}$ in the BCS–BEC crossover (dark blue line) for (a) Fermi gases and (b) lattice systems. Both cases display a dome-shaped behavior of $T_{\mathrm{c}}$ in the intermediate crossover regime but behave qualitatively different in the strong coupling BEC phase: $T_{\mathrm{c}}$ remains finite in the Fermi gas (a31 but approaches zero in the lattice case (b). During the crossover, $\xi_{\mathrm{p}}$ is reduced to the order of interparticle spacing $r$ (a & b) and lattice constant $a$ (b). For the lattice system, we contrast the evolution towards the multiorbital strong coupling phase (light blue line) discussed in this article. Here, the localization of pairs differently affects the bosonic hopping $t_{\mathrm{B}}$, as drawn in the insets on the right. In the lattice BEC limit, bosonic hopping relies on a second-order process involving two fermionic hoppings, $t_{\mathrm{F}}$, and a virtual intermediate state with broken pairs that is inhibited by the strong attraction $g<0$. Consequently, this process and also $T_{\mathrm{c}}\propto t_{\mathrm{B}}=t_{\mathrm{F}}^{2}/|g|$ are quenched at large $|g|$ 24, 21. In the multiorbital strong coupling case, the local coexistence of paired and unpaired electrons fluctuating between different orbitals enables bosonic hopping $t_{\mathrm{B}}$ as a first order process in $t_{\mathrm{F}}$ without any intermediate broken pair states. See Supplementary Note 7 for details on the involved states and energy diagram specific to the case of A3C60. A second temperature scale $T^{*}$ is drawn in both panels as its splitting from $T_{\mathrm{c}}$ (corresponding to the opening of a pseudogap) marks the beginning of the crossover regime 24.
@@
In this work, we demonstrate how multiorbital effects32, 33 can enhance superconductivity beyond the expectations of the lattice BCS–BEC crossover phenomenology (as contrasted in Fig. 1b) with a model inspired by alkali-doped fullerides (A3C60 with A=K, Rb, Cs). The material family of A3C60 hosts exotic $s$-wave superconductivity of critical temperatures up to $T_{\mathrm{c}}=38$ K, being the highest temperatures among molecular superconductors 34, 35, 36, 37, and they possibly reach photo-induced SC at even higher temperatures 15, 17, 18, 38. In order to theoretically characterize the SC state, the knowledge of the intrinsic SC length scales is essential. While BCS theory and Eliashberg theory provide a microscopic description of $\lambda_{\mathrm{L}}$ and $\xi_{0}$ for weakly correlated materials 39, 40, 41, their validity is unclear for superconductors with strong electron correlations. To the best of our knowledge, $\xi_{0}$ is generally not known from theory in strongly correlated materials. Only approaches to determine $\lambda_{\mathrm{L}}$ exist where an approximate, microscopic assessment of the SC stiffness from locally exact theories has been established, albeit neglecting vertex corrections 42, 43, 44, 45, 46, 47.
@@
For this reason, we introduce a novel theoretical framework to microscopically access $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from the tolerance of SC pairing to spatial OP variations. Central to our approach are calculations in the superconducting state under a constraint of finite-momentum pairing (FMP). Via Nambu–Gor’kov Green functions we get direct access to the superconducting OP and the depairing current $j_{\mathrm{dp}}$ which in turn yield $\xi_{0}$ and $\lambda_{\mathrm{L}}$. The FMP constraint is the SC analog to planar spin spirals applied to magnetic systems 48, 49, 50. As in magnetism, a generalized Bloch theorem holds that allows us to consider FMP without supercells; see Supplementary Note 2 for a proof. As a result, our approach can be easily embedded in microscopic theories and ab initio approaches to tackle material-realistic calculations.
@@
In this work, we implement FMP in Dynamical Mean-Field theory (DMFT) 51 to treat strongly correlated superconductivity. In DMFT, the interacting many-body problem is solved by self-consistently mapping the lattice model onto a local impurity problem. By this, local correlations are treated exactly. We apply the FMP-constrained DMFT to A3C60 where we find $\xi_{0}$ and $\lambda_{\mathrm{L}}$ in line with experiment for model parameter ranges derived from ab initio estimates, validating our approach. For enhanced pairing interaction, we then reveal a multiorbital strong coupling SC state with minimal $\xi_{0}$ on the order of only 2 – 3 lattice constants, but with robust stiffness $D_{\mathrm{s}}$ and high $T_{\mathrm{c}}$ which increases with the pair interaction strength without a dome shape. This strong coupling SC state is distinct to the lattice BCS–BEC crossover phenomenology showing promising routes to optimize superconductors with higher critical temperatures. We discuss this possibility in-depth after the presentation of our results.
The remaining paper is organized as follows: First, we motivate how the FMP constraint is linked to $\lambda_{\mathrm{L}}$ and $\xi_{0}$ from phenomenological Ginzburg–Landau theory after which we summarize the microscopic approach from Green function methods. Technical details are available in the Supplementary Information as cross-referenced at relevant points. Subsequently, we discuss our results for the multiorbital model of A3C60. Readers primarily interested in the analysis of these results might skip the first part on the physical motivation for obtaining superconducting length scales from the FMP constraint.
@@ Results
Extracting superconducting length scales from the constraint of finite-momentum pairing
@@
In most SC materials, Cooper pairs do not carry a finite center-of-mass momentum $\bm{q}=\bm{0}$. Yet, in presence of external fields, coexisting magnetism, or even spontaneously SC states with FMP, i.e., $\bm{q}\neq\bm{0}$, might arise 52, 53, 54, 55, 56 as originally conjectured in Fulde–Ferrel–Larkin–Ovchinnikov (FFLO) theory 57, 58, 59. Here, we introduce a method to access the characteristic SC length scales $\xi_{0}$ and $\lambda_{\mathrm{L}}$ of strongly correlated materials through the calculation of FMP states with a manually constrained pair center-of-mass momentum $\bm{q}$40, 4, 60, 61, 62. For this, we enforce the OP to be of the FMP form $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ corresponding to FF-type pairing 57, which is to be differentiated from pair density waves with amplitude modulations 58, 59, 63. We contrast the OP for zero and finite momentum in real and momentum space in the top panel of Fig. 2. For FMP, the OP’s phase is a helix winding around the direction of $\bm{q}$, while the OP for zero-momentum pairing is simply a constant.
@@
Before turning to the implementation in microscopic approaches, we motivate how the FMP constraint relates to $\lambda_{\mathrm{L}}$ and $\xi_{0}$. The Ginzburg–Landau (GL) framework provides an intuitive picture to this connection40, 4, 61 which we summarize here and discuss in detail in Supplementary Note 1, including a derivation of the following equations. The GL low-order expansion of the free energy density $f_{\mathrm{GL}}$ in terms of the FMP-constrained OP reads
$f_{\mathrm{GL}}[\Psi_{\bm{q}}]=\alpha|\Psi_{\bm{q}}|^{2}+\frac{b}{2}|\Psi_{\bm{q}}|^{4}+\frac{\hbar^{2}}{2m^{*}}q^{2}|\Psi_{\bm{q}}|^{2}$ (1)
@@
with $q=|\bm{q}|$. $\alpha$, $b$, and $m^{*}$ are the material and temperature dependent GL parameters. Here, the temperature-dependent (GL) correlation length $\xi(T)$ appears as the natural length scale of the amplitude mode ($\propto|\alpha|$) and the kinetic energy term
$\xi(T)=\sqrt{\frac{\hbar^{2}}{2m^{*}|\alpha|}}=\xi_{0}\left(1-\frac{T}{T_{\mathrm{c}}}\right)^{-\frac{1}{2}}$ (2)
@@
with its zero-temperature value $\xi_{0}$ being the coherence length 4. The stationary point of Eq. (1) shows that the $\bm{q}$-dependent OP amplitude $\left|\Psi_{\bm{q}}\right|=\left|\Psi_{0}\right|\sqrt{1-\xi^{2}q^{2}}$ ($T$-dependence suppressed) decreases with increasing momentum $q$. For large enough $q$, SC order breaks down (i.e., $\Psi_{\bm{q}}\to 0$) as the kinetic energy from phase modulations becomes comparable to the gain in energy from pairing. The length scale associated with this breakdown is $\xi(T)$ and can, therefore, be inferred from the $q$-dependent OP suppression. We employ, here, the criterion $\xi(T)=1/(\sqrt{2}|\bm{Q}|)$ with $\bm{Q}$ such that $|\Psi_{\bm{Q}}(T)/\Psi_{0}(T)|=1/\sqrt{2}$ for fixed $T$ (see Supplementary Note 5-A for more information).
@@
The finite center-of-mass momentum of the Cooper pairs entails a charge supercurrent $\bm{j}_{\bm{q}}\propto|\Psi_{\bm{q}}|^{2}\bm{q}$, c.f. Eq. (8) in Supplementary Note 1. This current density is a non-monotonous function of $q$ with a maximum called depairing current density $j_{\mathrm{dp}}$. It provides a theoretical upper bound to critical current densities, $j_{\mathrm{c}}$, measured in experiment. We note that careful design of SC samples is necessary for $j_{\mathrm{c}}$ reaching $j_{\mathrm{dp}}$ as its value crucially depends on sample geometry and defect densities 40, 61, 64. As the current is related to the London penetration depth $\lambda_{\mathrm{L}}$ through the second London equation $\mu_{0}\bm{j}=-\lambda_{\mathrm{L}}^{-2}\bm{A}$, we can determine $\lambda_{\mathrm{L}}$ from $j_{\mathrm{dp}}$ in GL theory via
$\lambda_{\mathrm{L}}(T)=\sqrt{\frac{\Phi_{0}}{3\sqrt{3}\pi\mu_{0}\xi(T)j_{\mathrm{dp}}(T)}}=\lambda_{\mathrm{L},0}\left(1-\left(\frac{T}{T_{\mathrm{c}}}\right)^{4}\right)^{-\frac{1}{2}}$ (3)
@@
with the magnetic flux quantum $\Phi_{0}=h/2e$. The temperature dependence with the quartic power stated here is empirical 39, 40 and we find that it describes our calculations better than the linearized GL expectation as discussed in Supplementary Note 5-B. Note that taking the $q\to 0$ limit in our approach is related to linear-response-based methods to calculate $\lambda_{\mathrm{L}}$ or, equivalently, the stiffness $D_{\mathrm{s}}$ 62, 47.
@@
Fig. 2: Influence of finite-momentum pairing (FMP) constraint on the superconducting condensate. The top panel insets sketch the position and momentum space representation of the order parameter (OP) $\Psi_{\bm{q}}(\bm{r})=|\Psi_{\bm{q}}|\mathrm{e}^{i\bm{qr}}$ in the zero-momentum (left, $q=0$) and finite-momentum pairing states (right, $q>0$). The main panels show (a) the momentum dependence of the OP modulus and (b) the supercurrent density $j_{q}=|\bm{j}_{\bm{q}}|$ as function of Cooper pair momentum $q=|\bm{q}|$ in reciprocal lattice units (r.l.u.). Gray lines indicate the points of extracting $\xi$ and $j_{\mathrm{dp}}$ (see text). The data shown are results for the A3C60 model (c.f. Eq. (7)) with lattice constant $a$ and interaction parameters $U/W=1.4$, $J/W=-0.04$ evaluated at different temperatures $T$ (color coded; see color bar with white triangular markers).
@@
The GL analysis shows that the OP suppression and supercurrent induced by the FMP constraint connect to $\xi_{0}$ and $\lambda_{\mathrm{L}}$. In a microscopic description, we acquire the OP and supercurrent density from the Nambu–Gor’kov Green function
@@
where $\psi_{\bm{k},\bm{q}}^{\dagger}=\left(c_{\bm{k}+\frac{\bm{q}}{2}\uparrow}^{\dagger}\ \ c_{-\bm{k}+\frac{\bm{q}}{2}\downarrow}\right)$ (orbital indices suppressed) are Nambu spinors that carry an additional dependence on $\bm{q}$ due to the FMP constraint. $G$ ($F$) denotes the normal (anomalous) Green function component for electrons ($G$) and holes ($\bar{G}$) in imaginary time $\tau$. For $s$-wave superconductivity as in A3C60 65, 66, 67, 68, we use the orbital-diagonal, local anomalous Green function as the OP
@@
which is the same for all orbitals $\alpha$. This allows us to work with a single-component OP. The current density can be calculated via (c.f. Eq. (26) in the Supplementary Information)